OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern?

OpenStudy (anonymous):

A. an = an – 1 + 3n – 2 B. an = an – 1 + 2n – 3 C. an = an – 1 + (3n – 2) D. an = an – 1 + (2n – 3)